A mass of gas has an initial pressure of 1 bar and a temperature of 20 0C. The temperature of the gas is now increased to 550 0C while the volume remains constant. Determine the final pressure of the gas.


1
Expert's answer
2021-10-10T09:34:50-0400

Q249210

A mass of gas has an initial pressure of 1 bar and a temperature of 20 0C. The temperature of the gas is now increased to 550 0C while the volume remains constant. Determine the final pressure of the gas.


Solution :


The initial pressure and temperature of the gas are 1 bar and 20 oC respectively.

Here volume is kept constant and the temperature is increased to 550 0C.


We can use the law which relates the pressure and temperature and keeping all other things like moles, and volume constant.


Amonton's law relates pressure and temperature at constant volume.


For two different sets of conditions, the law can be written as



"\frac{P_1}{T1} = \frac{P_2}{T_2}"


In the question, we are given the temperatures in degrees celsius, we will have to convert it to Kelvin.

P1 = 1 bar,

T1 = 20 0C = 20 + 273.15 = 293.15 K


P2 = unknown,

T2 = 550 0C = 550 + 273.15 = 823.15 K


plug all this information in the formula we have



"\frac{1 \ bar }{293.15 \ K } = \frac{P_2}{823.15 \ K }"

Arranging this equation for P2, we have



"\frac{1 \ bar }{293.15 \ K } * 823.15 \ K = P_2"


"2.8079 \ bar = P_2"


In the correct significant figure, the answer is 3 bar.


Hence the final pressure of the gas will be 3 bar.